OpenStudy (anonymous):

The value of an uncirculated mint state -65' 1950 Jefferson nickel minted in Denver is 5/3 the value of a 1945 nickel minted in Philadelphia in similar conditions. Together the total value of the two coins is $56. What is the value of each coin? 1945 nickel =_______ 1950 nickel=________

OpenStudy (anonymous):

Let the value of the 1945 nickel to be x. Then (5/3)x + x = 56 (8/3)x = 56 x = 21 So 1945 nickel = $21, 1950 nickel = 35